Intercontinental, Inc., provides you with the following data for its single product: Sales price per unit Fixed costs (per month) 52.00 Selling, general, and administrative (SG&A) Manufacturing overhead 1,350,000 2,700,000 Variable costs (per unit) Direct labor Direct materials Manufacturing overhead SG&A 7.00 12.00 10.00 5.00 Number of units produced per month 300,000 units Required Compute the amounts for each of the following assuming that both production levels are within the relevant range. (Do not round intermediate calculations. Round your answers to 2 decimal places.) 300,000 units 400,000 units a. Prime cost per unit. b. Contribution margin per unit. C. Gross margin per unit. d. Conversion cost per unit. e. Variable cost per unit. f. Full absorption cost per unit. g. Variable production cost per unit. h. Full cost per unit.
Answer #1
300000 units($) 400000 units($)
Prime cost per unit 19 19
Contribution margin per unit 18 18
Gross margin per unit 14 16.25
Conversion cost per unit 26 23.75
Variable cost per unit 34 34
Full absorption cost per unit 38 35.75
Variable production cost per unit 29(7+12+10) 29(7+12+10)

Prime cost per unit= direct labour per unit + direct material per unit

= 7+ 12=$19

Total varible cost per unit= direct labour per unit+direct material per unit+ variable manufacturing overhead + sg&a per unit

= 7+12+10+ 5= $34

Contribution margin per unit= selling price per unit - variable cost per unit = 52- 34 = $18

Note Since variable cost per unit does not vary with change in number of units, contribution per unit and prime cost per unit will be same for both production levels.

Cost of goods sold per unit= direct labour per unit+ direct material per unit+ variable manufacturing overhead per unit+ fixed manufacturing overhead per unit

  1. For 300000units= 7+12+10+9 (2700000/300000=9)= $38
  2. For 400000units= 7+12+10+6.75 (2700000/400000=6.75)= $35.75

Gross profit margin per unit= selling price per unit- cost of goods sold out unit

  1. For 300000units= 52-38= $14
  2. For 400000units= 52-35.75= $16.25

Conversion cost per unit= direct labour per unit + variable manufacturing overhead per unit+ fixed manufacturing overhead per unit

  1. For 300000units = 7+10+9= $26
  2. For 400000 units= 7+10+ 6.75=$23.75